# Knowledge Entropy Decay during Language Model Pretraining Hinders New Knowledge Acquisition

### v1

@@ B.2 Metric
Given a language model $\theta$, $\theta_{\text{PT}}$ represents the model extracted from a pretraining step and serves as the initial point for continual learning and $\theta_{\text{CL}}$ represents the model after it. The acquisition metric $\mathcal{A}(\theta)$ for the model $\theta$ is defined as Equation 4. When given a corpus set of once setting $C_{\text{once}}$, each instance in a corpus $c_{i}$ has a corresponding set of probes $P_{c_{i}}$, which contains 15 different probes. To calculate the performance of the once setting, $\mathcal{K}_{\text{once}}(\theta)$, we compute the average log probability $\ell$($p_{i}$;$\theta$) of target span for each probe $p_{i}\in P_{c_{i}}$ across all instances in $c_{i}\in C_{\text{once}}$ and sum these averages. The same calculation is performed for the paraphrase setting. The total performance $\mathcal{K}(\theta)$ is calculated by the weighted average of $\mathcal{K}_{\text{once}}(\theta)$ and $\mathcal{K}_{\text{para}}(\theta)$. Finally, the acquisition metric $\mathcal{A}(\theta)$ is defined as the improvement rate in performance $\mathcal{K}(\theta)$ from the initial model state $\theta_{\text{PT}}$ to final model state $\theta_{\text{CL}}$.
$\mathcal{K}_{\text{once}}(\theta)=\frac{1}{|C_{\text{once}}|}\sum_{c_{i}\in C_{\text{once}}}\frac{1}{|P_{c_{i}}|}\sum_{p_{i}\in P_{c_{i}}}\ell(p_{i};\theta);\qquad\mathcal{K}_{\text{para}}(\theta)=\frac{1}{|C_{\text{para}}|}\sum_{c_{i}\in C_{\text{para}}}\frac{1}{|P_{c_{i}}|}\sum_{p_{i}\in P_{c_{i}}}\ell(p_{i};\theta)$ (4)
$\mathcal{K}(\theta)=\frac{|C_{\text{once}}|\times\mathcal{K}_{\text{once}}(\theta)+|C_{\text{para}}|\times\mathcal{K}_{\text{para}}(\theta)}{|C_{\text{once}}|+|C_{\text{para}}|};\qquad\mathcal{A}(\theta)=\frac{\mathcal{K}(\theta_{\text{CL}})-\mathcal{K}(\theta_{\text{PT}})}{\mathcal{K}(\theta_{\text{PT}})}$
@@
The forgetting metric $\mathcal{F}(\theta)$ is calculated by the average performance degradation from the initial model $\theta_{\text{PT}}$ to the final model $\theta_{\text{CL}}$ in six downstream tasks $\mathcal{T}$: SciQ (Welbl et al., 2017), Winograde (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), OBQA (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019), and ARC Easy (Clark et al., 2018).
@@
Figure 10: Rates of knowledge acquisition $\mathcal{A}(\theta)$ and final performance of knowledge probe $\mathcal{K}(\theta)$ according to the number of injection. Paraphrase prompt (solid line) were injected 10 times and Once (dotted line) prompt only once throughout continual training.
@@ B.3 Frequency of Knowledge Injections
We divide the experiment into two settings: the once setting, where knowledge is injected a single time, and the paraphrase setting, where knowledge is injected ten times using ten paraphrased paragraphs. Figure 11 shows knowledge acquisition results based on the frequency of injections. Knowledge acquisition and final performance generally follow similar trends in both settings, with models in the later stages of pretraining showing the lowest performance. However, the performance and acquisition rate of once setting lags behind that of paraphrase setting. Also, notably, for models in the final stage of pretraining, the acquisition rate in the once setting was negative. This indicates that the log probability of the injected knowledge decreased, preventing successful incorporation of the new knowledge. In other words, even the knowledge injected during continual learning is subject to forgetting throughout the continual learning process.
@@ B.4.1 Baseline Setup
Our base experiements are reported with hyperparameter configuration most closely aligned with continual knowledge learning studies, specifically with batch size 128, learning rate 4e-4, while training single-epoch of PubMed corpus. We use adamW optimizer ($\beta$ = 0.9, 0.95, weight decay= 0.1), cosine LR scheduler with warmup=0.05, and set maximum sequence length as 1024. We randomly selected 204,800 instances from the PubMed and C4 datasets, and matched the sequence length to 1,024 tokens by concatenating instances. This resulted in a training dataset consisting of roughly 210 million tokens.
As analyzed in 4.2, final performance generally deteriorates as later-stage models were utilized as initial model. Figure 11 illustrates the model’s initial performance before continual learning, as well as its performance afterward. Models in the later stages of pretraining exhibit superior language modeling abilities before continual learning, as evidenced by the lower log probability for the newly injected knowledge (dotted line). However, after continual learning, their performance deteriorates compared to models from the earlier stages of pretraining (solid line). Similarly, the downstream task performance of the later-stage models was better initially, but after continual learning, their performance declined more than that of the earlier-stage models.
@@ B.4.2 Various Settings
We observed consistently, even in settings where hyperparameters were altered, that models in the later stages of pretraining struggle to learn new knowledge and retain existing knowledge.
@@
To test universal deterioration of knowledge acquisition and retention capabilities, we also experimented with OLMo 7B model Groeneveld et al. (2024). In Figure 4, the 7B model exhibits a clear trend of diminishing knowledge acquisition $\mathcal{A}(\theta)$ capabilities as pre-training progresses. This decline is accompanied by an increase in forgetting $\mathcal{F}(\theta)$, indicating that the model struggles to retain previously learned information as new data is introduced.
@@
In Table 2 line (b), when the batch size is large, the influence of individual data points on the model update decreases, resulting in a lower acquisition rate, while forgetting is less pronounced. In the later stages of training, however, if sufficient learning rate warmup is not provided, the model seems to collapse.
@@
In Table 2 line (c), when the learning rate is increased, not only is the acquisition rate lower, but forgetting becomes more pronounced. However, when the learning rate is very small (in line (d)), the models in the late stage shows improved acquisition performance with least gap compared with initial models, but still remains inferior compared to the mid stage model.
@@
B.5 Resuscitation Experiment: Varying $p$ while Fixing $q$
Figure 13 shows the overall performance when $q$ was fixed at 1, meaning that the lowest $p$of coefficients were multiplied to converge toward the layer’s mean. It is shown that increasing $p$ to activate a larger portion of inactive parameters generally led to improved performance. However, interestingly, when $p$ was set too high, at 90, this resulted in a negative impact on performance, likely due to the unintended effect of reducing the parameters in already active regions.
@@ B.5 Resuscitation Experiment: Varying $p$ while Fixing $q$
B.6 Resuscitation Experiment across pretraining steps
Figure 14 illustrates the overall performance when $q$ was fixed to 2 and $p$ to 50, across different pretraining stages of the original model. The effect of resuscitation method becomes more pronounced as the original model progresses to later stages of pretraining, as indicated by the transition from the dotted line to the solid line. We hypothesize that this is because late-stage models tend to rely on a smaller subset of memory sources and thus benefit from a broader scope of activation enabled by the resuscitation method. As shown in Figure 14, end performance deteriorates when the beginning model is initial (118k) and mid (369k) stage model, indicating that resuscitation may impair performance when the model’s knowledge entropy is not low enough. This trend of the resuscitation showing a more positive effect for models in later stage of pretraining can also be seen in Figure 15, which shows the result when varying $q$ while fixing $p$ as 50: performance deteriorates when running continual learning on model from 369k, while improvement of performance with larger $q$ is observed when model is from 554k.

### v3

@@ B.2 Metric
Given a language model $\theta$, $\theta_{\text{PT}}$ represents the model extracted from a pretraining step and serves as the initial point for continual learning, and $\theta_{\text{CL}}$ represents the model after it. The acquisition metric $\mathcal{A}(\theta)$ for the model $\theta$ is defined as Equation 4. When given a corpus set of once setting $W_{\text{once}}$, each instance in a corpus $w_{i}$ has a corresponding set of probes $P_{w_{i}}$, which contains 15 different probes. To calculate the performance of the once setting, $\mathcal{K}_{\text{once}}(\theta)$, we compute the average log probability $\ell$($p_{i}$;$\theta$) of the target span for each probe $p_{i}\in P_{w_{i}}$ across all instances in $w_{i}\in W_{\text{once}}$ and sum these averages. The same calculation is performed for the paraphrase setting. The total performance $\mathcal{K}(\theta)$ is calculated by the weighted average of $\mathcal{K}_{\text{once}}(\theta)$ and $\mathcal{K}_{\text{para}}(\theta)$. Finally, the acquisition metric $\mathcal{A}(\theta)$ is defined as the improvement rate in performance $\mathcal{K}(\theta)$ from the initial model state $\theta_{\text{PT}}$ to final model state $\theta_{\text{CL}}$.
$\mathcal{K}_{\text{once}}(\theta)=\frac{1}{|W_{\text{once}}|}\sum_{w_{i}\in W_{\text{once}}}\frac{1}{|P_{w_{i}}|}\sum_{p_{i}\in P_{w_{i}}}\ell(p_{i};\theta);\qquad\mathcal{K}_{\text{para}}(\theta)=\frac{1}{|W_{\text{para}}|}\sum_{w_{i}\in W_{\text{para}}}\frac{1}{|P_{w_{i}}|}\sum_{p_{i}\in P_{w_{i}}}\ell(p_{i};\theta)$ (4)
$\mathcal{K}(\theta)=\frac{|W_{\text{once}}|\times\mathcal{K}_{\text{once}}(\theta)+|W_{\text{para}}|\times\mathcal{K}_{\text{para}}(\theta)}{|W_{\text{once}}|+|W_{\text{para}}|};\qquad\mathcal{A}(\theta)=\frac{\mathcal{K}(\theta_{\text{CL}})-\mathcal{K}(\theta_{\text{PT}})}{\mathcal{K}(\theta_{\text{PT}})}$
@@
The forgetting metric $\mathcal{F}(\theta)$ is calculated as the average performance degradation from the initial model $\theta_{\text{PT}}$ to the final model $\theta_{\text{CL}}$ across six downstream tasks $\mathcal{T}$: SciQ (Welbl et al., 2017), Winograde (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), OBQA (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019), and ARC Easy (Clark et al., 2018).
@@
Figure 11: Rates of knowledge acquisition $\mathcal{A}(\theta)$ and final performance of knowledge probe $\mathcal{K}(\theta)$ according to the number of injection. Paraphrase prompts (solid line) were injected 10 times and Once (dotted line) prompt only once throughout continual training.
@@ B.3 Frequency of Knowledge Injections
We divide the experiment into two settings: the once setting, where knowledge is injected a single time, and the paraphrase setting, where knowledge is injected ten times using ten paraphrased paragraphs. Figure 12 presents knowledge acquisition results based on the frequency of injections. Knowledge acquisition and final performance generally follow similar trends in both settings, with models in the later stages of pretraining showing the lowest performance. However, the performance and acquisition rate of once setting lag behind those of paraphrase setting. Also, notably, for models in the final stage of pretraining, the acquisition rate in the once setting was negative. This indicates that the log probability of the injected knowledge decreased, preventing successful incorporation of the new knowledge. In other words, even the knowledge injected during continual learning is subject to forgetting throughout the continual learning process.
@@ B.4.1 Baseline Setup
Our base experiments are conducted using a hyperparameter configuration most closely aligned with continual knowledge learning studies, specifically with a batch size of 128, a learning rate of 4e-4, while training single-epoch of PubMed corpus. We use adamW optimizer ($\beta$ = 0.9, 0.95, weight decay= 0.1), cosine LR scheduler with warmup=0.05, and set maximum sequence length as 1024. We randomly selected 204,800 instances from the PubMed and C4 datasets and matched the sequence length to 1,024 tokens by concatenating instances. This resulted in a training dataset consisting of approximately 210 million tokens.
As analyzed in Section 4.2, final performance generally deteriorates as later-stage models are utilized as the initial model. Figure 12 illustrates the model’s initial performance before continual learning, as well as its performance afterward. Models in the later stages of pretraining exhibit superior language modeling abilities before continual learning, as evidenced by the lower log probability for the newly injected knowledge (dotted line). However, after continual learning, their performance deteriorates compared to models from the earlier stages of pretraining (solid line). Similarly, the downstream task performance of the later-stage models was better initially, but after continual learning, their performance declined more than that of the earlier-stage models.
@@ B.4.2 Various Settings
We observed consistently, even with altered hyperparameter settings, that models in the later stages of pretraining struggle to learn new knowledge and retain existing knowledge.
@@
To test the universal deterioration of knowledge acquisition and retention capabilities, we also experimented with the OLMo 7B model Groeneveld et al. (2024). In Figure 4, the 7B model exhibits a clear trend of diminishing knowledge acquisition $\mathcal{A}(\theta)$ capabilities as pretraining progresses. This decline is accompanied by an increase in forgetting $\mathcal{F}(\theta)$, indicating that the model struggles to retain previously learned information as new data is introduced.
@@
In Table 1 line (b), when the batch size is large, the influence of individual data points on the model update decreases, resulting in a lower acquisition rate, while forgetting is less pronounced. In the later stages of training, however, if sufficient learning rate warmup is not provided, the model appears to collapse.
@@
In Table 1 line (c), when the learning rate is increased, not only is the acquisition rate lower but forgetting becomes more pronounced. Conversely, when the learning rate is very small (in line (d)), the models in the late stage show improved acquisition performance with the least gap compared with initial models, but still perform worse than the mid-stage model.
@@
B.5 Training Dynamics of Pythia
Figure 14 shows the training dynamics of Pythia Biderman et al. (2023) 1.4B and OLMo 1B. We observe that the overall training dynamics of Pythia and OLMo are similar. However, Pythia tends to conclude early in training, using only 10% of training tokens compared to OLMo. As a result, Pythia struggles to capture the full dynamics of the language model during pretraining. This is because models in the early stage of training are still influenced by a large portion of randomly initialized parameters, which leads to different behavior compared to models that have undergone more training steps and reached a stable state. For OLMo, based on our analysis, we assume the model reaches a stable point around 118k steps (15% of the training dataset). However, Pythia’s final step occurs before this stable point, causing its trend to resemble that of OLMo but deviate from the trend observed in Section 4, where knowledge acquisition improves until the stable point and then decreases. We hypothesize that previous studies using intermediate checkpoints during pretraining (Chang et al., 2024; Sun & Dredze, 2024) also only measured with OLMo without Pythia for the same reason.
Figure 14: Knowledge Entropy and Rates of knowledge acquisition $\mathcal{A}(\theta)$ of OLMo 1B and Pythia 1.4B model across different pretraining stages
B.6 Resuscitation Experiment: Varying $p$ while Fixing $q$
@@
Figure 15 shows the overall performance when $q$ was fixed at 1, meaning that the lowest $p$of coefficients were scaled to converge toward the layer’s mean. It is shown that increasing $p$ to activate a larger portion of inactive parameters generally led to improved performance. However, interestingly, setting $p$ too high(e.g., at 90) negatively impacted performance, likely due to the unintended effect of reducing the parameters in already active regions.
@@
B.7 Resuscitation Experiment across Pretraining Steps
Figure 16 illustrates the overall performance when $q$ was fixed to 2 and $p$ to 50, across different pretraining stages of the original model. The effect of the resuscitation method becomes more pronounced as the original model progresses to later stages of pretraining, as indicated by the transition from the dotted line to the solid line. We hypothesize that this is because late-stage models tend to rely on a smaller subset of memory sources and thus benefit from a broader scope of activation enabled by the resuscitation method. As shown in Figure 16, end performance deteriorates when the beginning model is initial (118k) and mid (369k) stage model, indicating that resuscitation may impair performance when the model’s knowledge entropy is not sufficiently low. This trend of the resuscitation showing a more positive effect for models in the later stage of pretraining can also be seen in Figure 17, which shows the results when varying $q$ while fixing $p$ at 50: performance deteriorates when running continual learning on the model from 369k, while improvement of performance with larger $q$ is observed when the model is from 554k.
B.8 Resuscitation Experiment across Attention Layers
Figure 18: (a) Rates of knowledge acquisition $\mathcal{A}(\theta)$ and forgetting $\mathcal{F}(\theta)$ (b) End performance of knowledge probe $\mathcal{K}(\theta)$ and downstream task $\mathcal{P}(\theta)$ when resuscitation is applied at attention layers, with the x-axis representing temperature. The dotted line indicates performance when resuscitation is applied to the feed-forward layer with $p$=50, $q$=2. The dashed line represents the original performance without any resuscitation.
In Figure 3, we observe that as pretraining progresses, not only knowledge entropy but also attention entropy decreases. We explore whether extending the resuscitation method to the attention layers could similarly improve knowledge acquisition and retention. For the feed-forward layer, we can adjust knowledge entropy by tuning the parameters at positions with low activation values. However, since attention entropy reflects the probability distribution over token positions, the same resuscitation method cannot be directly applied to the attention layers. To artificially increase attention entropy, we employed temperature scaling on the softmax function in the attention calculation, which reduces the sparsity of the probability distribution.
We experimented with temperature values ranging from 1.5 to 3.0. Figure 18 shows that resuscitation in the feed-forward layer consistently results in the highest knowledge acquisition rate and generally leads to the best knowledge retention. This trend is observed in both knowledge probe results and downstream performance. These findings suggest that knowledge entropy plays a critical role in influencing both knowledge acquisition and forgetting, thereby driving the observed differences in performance.